OpenStudy (anonymous):

a line that contains (4,-4)and (-2,0)

OpenStudy (amistre64):

\[y = \frac{-4-0}{4+2}(x-4)-4\]

OpenStudy (anonymous):

(y-0)/(x+2)=(0+4)/(-2-4) =>-6y=4x+8

OpenStudy (anonymous):

What is your question ?

OpenStudy (amistre64):

..... missed it; 2*4 aint 6 :)

OpenStudy (anonymous):

y=-2x/3-3/2